OpenStudy (anonymous):

I need to solve Log base 5 of 5times the squareroot of 5

OpenStudy (unklerhaukus):

\[\log_5\big(5\sqrt 5\big)\]

OpenStudy (unklerhaukus):

\[\log_a(xy)=\log_a(x)+\log_b(y)\]

OpenStudy (anonymous):

3/2

OpenStudy (unklerhaukus):

\[\log_a(a)=1\] \[\log_a(z^n)=n\log_a(z)\]

OpenStudy (anonymous):

yes, thank you, but i'm not quite sure how to solve in exponential form. The answer I know is 3/2 but how?

OpenStudy (unklerhaukus):

\[\log_5\big(5\sqrt 5\big)=\log_5(5)+\log_5(\sqrt 5)=\log_5(5)+\log_5(5^{1/2})=\log_5(5)+\frac12\log_5(5)\]

OpenStudy (anonymous):